OpenStudy (anonymous):

triangle abc has verticies a(0,0) b(-4,0)and c(-2,4). the coordinates of each vertex are multiplied by 3 to create triangle aef determin if triangle abc is similar to triangle aef. explain your answer.

OpenStudy (imer):

Yes they are similar because the length is increased by factor of "3" and and area is increased by factor of "3^2".

OpenStudy (imer):

Area of original triangle ABC Triangle,\[\frac{ 1 }{ 2 }*4*4=8\] and Area of Triangle AEF \[\frac{ 1 }{ 2 }*12*12=72\] And therefore, the ratio is \[\frac{ 72 }{ 8 }=9=3^2\]